Jj0002:

Evaluate the expression \((4x^3y^{-2})(3x^{-2}y^4)\)

jhonyy9:

please post this jpg. file using the Attach File button below to this window

jhonyy9:

use this property a^x *a^y = a^(x +y) and a^x : a^y = a^(x - y )

jhonyy9:

ok step by step 4*3 = ?

jhonyy9:

bravo who asked is offline

jhonyy9:

@Laylalyssa

Laylalyssa:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 ok step by step 4*3 = ? \(\color{#0cbb34}{\text{End of Quote}}\) 12

jhonyy9:

ok the second terms x^3 *x^(-2) = ?

jhonyy9:

use this formula what i ve wrote above

jhonyy9:

when you multiple the same terms with exponents so the exponents you need assume them

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 ok the second terms x^3 *x^(-2) = ? = x^(3-2) = x^? \(\color{#0cbb34}{\text{End of Quote}}\)

jhonyy9:

and in case of the next term of y just so in this way

jhonyy9:

@supie

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 ok the second terms x^3 *x^(-2) = ? = x^(3-2) = x^? \(\color{#0cbb34}{\text{End of Quote}}\) \(\color{#0cbb34}{\text{End of Quote}}\) so i just solve \(x^3 *x^(-2) = ? = x^(3-2)\)?

supie:

\(x^3 *x^{(-2)} = ? = x^{(3-2)}\)?